# Simulacra as Conscious Exotica

### v1

@@ 1   Introduction
As the behaviour of generative AI systems, especially conversational agents based on large language models (LLMs), becomes more compellingly human-like, so the temptation to ascribe human qualities to them will become increasingly hard to resist. The temptation will be especially great if those agents are virtually embodied, if they interact with users in immersive virtual worlds through avatars with a human- or animal-like form. [^1] The term “agent” is used liberally throughout this paper. Although it has anthropomorphic overtones, it is used as a technical term in AI (Russell and Norvig, 2010, Chapter 2) that doesn’t imply agency in its fullest sense (Schlosser, 2019). Consciousness is one such quality, one that is especially morally valenced. To see a fellow creature as conscious (or sentient) often goes hand-in-hand with the sense that we should behave decently towards it (Singer, 1975; Nussbaum, 2023). To see an AI system as conscious would be to admit it into this fellowship of conscious beings, and potentially to give it “moral standing”, which would be a serious matter (Ladak, 2023).
Unfortunately, the ascription of consciousness, except in the ordinary case of human beings, is a tricky business, and the less human-like the candidate, the more problematic the ascription. [^2] Even in humans, there are difficult clinical cases. Yet philosophical intuition is inclined to grant the possibility that, within the space of possible minds – a space that encompasses not only humans and other animals, but also putative extraterrestrial life-forms as well as sufficiently advanced kinds of artificial intelligence (Sloman, 1984) – there exist conscious exotica, entities that are extremely different from anything found in (terrestrial) biology but that belong to the fellowship of conscious beings (Nagel, 1974; Shanahan, 2016). [^3] The philosophical intuition in question manifests prominently in the tropes of science fiction, but this is no reason to sideline it (Rennick, 2021).
@@
The central concern of this paper is whether (virtually embodied) LLM-based conversational agents, despite being “mere” simulacra of human behaviour, could ever qualify as conscious exotica. [^4] Butlin et al. (2023) address a similar question using a very different methodology. See also (Chalmers, 2023). Following the treatment in (Shanahan, 2016), which draws heavily on the later work of Wittgenstein, the paper foregrounds embodied interaction as a basis for the use of consciousness language.
@@ 2   Language Models and AI
The core component of a contemporary conversational agent, such as OpenAI’s ChatGPT or Google’s Gemini, is a large language model, such as GPT-4 (OpenAI, 2023) or Gemini Ultra (Anil et al., 2023). The basic task of the LLM is to (probabilistically) generate continuations of sequences of words (tokens) (Vaswani et al., 2017; Shanahan, 2024). The LLM embedded in a typical conversational agent is the product of two steps. First, a base model is trained to perform next token prediction on a large corpus of textual data. Second, the base model is fine-tuned a) to be effective at following instructions in a dialogue setting, and b) to take account of feedback from humans raters with respect to toxicity, bias, and so on. The resulting LLM is embedded in a dialogue system that takes turns with the user to build up a conversation in the context of an initial dialogue prompt, not seen by the user, that sets the tone of the exchange.
@@
Today’s conversational agents are hardly without their limitations. They have a tendency to generate inaccurate, made-up information (a phenomenon often (mis-)termed “hallucination”), and their reasoning skills are poor. Nevertheless, the experience of interacting with them is sufficiently compelling, and their conversational capabilities are sufficiently close to human level, that the urge to speak of them in anthropomorphic terms is almost overwhelming.
Anthropomorphising AI systems is sometimes harmless fun, if the user is in the know, and sometimes useful for explaining and predicting the a system’s behaviour, as when we adopt what Dennett calls the intentional stance, the strategy of “interpreting the behavior of an entity … as if it were a rational agent” (Dennett, 2009). Anthropomorphism is problematic when it involves the misleading attribution of human properties to systems that lack those properties, giving rise to false expectations for how the system will behave.
@@ 3   Anthropomorphism and Role Play
The position adopted in this paper, following (Shanahan, 2024), is that we should be wary of taking too seriously talk of beliefs in simple LLM-based conversational agents, despite their impressive conversational skills, since they lack the means to “participate fully in the human language game of truth”. Specifically, a simple LLM-based conversational agent cannot measure its words against external reality and update what it says accordingly, a capacity that is central to the concept of belief in its fullest sense.
While it may not be appropriate to think of simple conversational agents as literally having beliefs, they can still usefully be thought of as role-playing or simulating an entity with beliefs (Andreas, 2022; Janus, 2022; Shanahan et al., 2023). LLMs encode a great deal of human knowledge, and a suitably fine-tuned and prompted base model will effectively play the part of a helpful assistant in a turn-taking setting, answering factual questions (more or less accurately) as if it believed, and had good reasons to believe, its own answers. In general, the role play framing allows us to use familiar folk-psychological terms to describe, explain, and predict the behaviour of LLM-based systems without falling into the trap of anthropomorphism.
@@
Now, to what extent can we extend this treatment of belief to other mental attributes, such as desires, goals, intentions, and, most pertinently, to consciousness? To a degree, the same trends that legitimise talk of belief in enhanced conversational agents apply to the concepts of goals and intentions. A conversational agent capable of deliberation and tool-use can play the part of an assistant that forms plans on behalf of the user, and sets about executing them by carrying out real-world actions such as making purchases, sending emails, and so on. In this narrow context, there is little to distinguish role-play from authenticity. Real emails are sent, real items are purchased. Thus the agent fulfils its intentions and meets its goals, hopefully to the satisfaction of a real user.
@@ 4   Wittgenstein Versus Dualism
The aim of this section is to present a philosophical approach that can be applied to exotic candidates for admission to the fellowship of conscious beings, such as LLM-based conversational agents. First, though, some remarks in lieu of a definition are in order. Consciousness is a multi-faceted concept. In everyday conversation, we speak of wakefulness, awareness, attention, experience, sensation, feeling, emotion, and so on. The scientific and philosophical literature supplements common speech with a distinctive vocabulary of its own: perception, introspection, phenomenology, sentience, selfhood, higher-order states, mental imagery, inner speech, and so on. [^5] This list deliberately omits the influential distinction between phenomenal consciousness and access consciousness introduced by Block (1995) in a related context, a distinction that is inimical to the present philosophical project. To speak generically of consciousness is to allude to this whole cloud of concepts, and an entity has the capacity for consciousness to the extent that this vocabulary is applicable to it.
@@
To achieve this dissolution we have to take on board Wittgenstein’s overarching philosophical project, wherein his view of language plays a central role. According to this view, language is an inherently embodied and social phenomenon, an aspect of human collective activity. So rather than asking what a word means, we should instead ask how it is used, what its role is in everyday human affairs. This applies no less to tricky philosophical words, such as “consciousness” and its relatives, than it does to everyday words like “flower” or “hello”. Philosophical puzzles arise when “language goes on holiday”, when philosophically difficult words are taken far from their natural home in everyday life and used in peculiar ways to “bewitch our intelligence”.
To undo the spell of dualism, in the case of words like “sensation”, “experience”, and “feeling”, is no easy matter, but in the private language remarks, Wittgenstein takes us through a series of steps with the aim of doing this. He shows that a word that purported to denote a purely private sensation could not have any possible use in our language. Only words whose correct usage can be adjudicated through what is public, by the community of language users, can have meaning. Wittgenstein is careful neither to deny nor to affirm the existence of the allegedly private, hidden thing, the “sensation itself”, so to speak. It is “not a something, but not a nothing either”. The conclusion, rather, is that “a nothing would serve as well as a something about which nothing can be said”.
@@ 5.3 A society-wide conversation
There is no guarantee of consensus here. Disagreement and debate are part of the conversation. [^8] Humphrey, for example, takes issue with the current (near) consensus on octopus sentience on the grounds that “they are not natural psychologists, they do not regard each other as selves, nor do they care” (Humphrey, 2022, p.205). Nor, as the conversation progresses, is there any guarantee of eventual convergence. Even if we do begin to treat an exotic entity as a conscious being and to describe it in such terms, as time goes by and more is learned about it, either at the level of behaviour or at the mechanistic level, this tendency might fade. Perhaps we will decide, collectively, that the language of consciousness is not the right one after all. Perhaps a little more nuance will be required. Perhaps a whole new vocabulary will emerge.
@@ 6.1 Simple conversational agents
Even simple conversational agents built on large language models, agents that that do nothing more than engage in textual dialogue, can elicit the feeling of a presence at the other end of the conversation (Schwitzgebel, 2023). So it is perhaps unsurprising that some users who have spent lots of time interacting with these agents will start to think of them as fellow conscious beings and to speak of them in such terms (Tiku, 2022; Colombatto and Fleming, 2023). [^9] If a human ascribes consciousness to an AI system even though they know that it is an artefact, then it passes the Garland Test, named after the 2015 film Ex Machina, written and directed by Alex Garland, in which the robot Ava is subjected to just such a test (Shanahan, 2016; Seth, 2021). The Garland Test is neutral about whether or not the artefact in question is really conscious (whatever that may mean). But how seriously should we take this?
According to the prescription of the previous section, for an exotic entity to qualify as as a candidate for the fellowship of conscious beings it must be possible to engineer an encounter with it, at least in principle if not in practice. However, it is not possible to engineer an encounter with a simple conversational agent, even in principle. This is because simple LLM-based conversational agents are not embodied; we cannot be with them in a shared world.
@@
To use the language of consciousness in a disembodied setting is to stray impossibly far from all this, its original home. If some community of language users insists on doing so, then one of two things must hold. Either they have bent the language of consciousness so radically out of shape that it has detached from its original nexus of meaning. Or, to the extent they believe this is not the case, that “consciousness” means the same for them as it always did for everyone else, they are clinging to the (alluring) dualistic picture of consciousness as a metaphysical kind whose very nature (private, hidden) means that it does not require embodiment in a world shared with others. [^10] There is a standard set of supposed counter-examples to this view that includes brains in vats, minds uploaded into computers, patients with locked-in syndrome, and perfectly normal people in sensory deprivation tanks. Each of these example is sometimes alleged to support the view that there can be consciousness without embodiment. However, in every case it is possible to engineer an encounter in the current sense. The brain in a vat can be re-connected to its body, the locked-in patient can be cured, the uploaded mind can be downloaded again, and the occupant of the sensory deprivation tank can be dragged back into the daylight.
@@ 6.2 Physical embodiment
The LLM-based generative AI systems deployed by today’s major tech corporations, such as OpenAI’s ChatGPT and Google’s Gemini, are more than simple conversational agents. Although language models remain their core component, they are multi-modal, tool-using conversational agents. As well as text, they can take images as input and generate images as output, and they can make calls to external apps (tools) such as calendars, calculators, search engines, and Python interpreters. Additionally, the architecture of these agents can be elaborated in ways that bring them closer to humans and animals, equipping them with a persistent memory of their interactions with the user and the world, for example, and by giving them the ability to plan and to form explicit lists of tasks and sub-goals (Xi et al., 2023).
@@
Is there any way to extend an LLM-based conversational agent into an AI system with which we could engineer an encounter, and that would thereby qualify as a candidate for consciousness? Indeed there is. We could embody it. One possibility is embodiment in physical robot form (Driess et al., 2023; Brohan et al., 2023). An LLM with multi-modal, tool-using capabilities is half-way there. Robot actions become another sort of tool, while data from the robot’s sensors (including camera images) is assimilated into its multi-modal input space. The LLM can then be suitably fine-tuned and prompted to carry out tasks for a user.
@@
Biological brains evolved to enable animals to survive and reproduce in complex environments, and language evolved to serve those fundamental needs. [^11] For some authors, consciousness is similarly bound up with these biological fundamentals. Seth, for example, writes: “[A]ll of our experiences and perceptions stem from our nature as self-sustaining living machines that care about their own persistence” (2021, p.255). In nature, therefore, embodiment is given, and language is inherently grounded in interaction with the physical world. In a robot controlled by an LLM, all this is back-to-front. A pre-trained statistical model of human language is bolted post hoc onto a robot body with no fundamental biological needs. The exoticism of the language model itself is compounded accordingly; a robot controlled by an LLM that exhibited human-like behaviour would be an especially exotic artefact.

### v2

@@ 1   Introduction
As the behaviour of generative AI systems, especially conversational agents based on large language models (LLMs), becomes more compellingly human-like, so the temptation to ascribe human qualities to them will become increasingly hard to resist. The temptation will be especially great if those agents are virtually embodied, if they interact with users in immersive virtual worlds through avatars with a human- or animal-like form. [^1] The term “agent” is used liberally throughout this paper. Although it has anthropomorphic overtones, it is used as a technical term in AI (Russell and Norvig, 2010, Chapter 2) that doesn’t imply agency in its fullest sense (Schlosser, 2019). Consciousness is one such quality, one that is especially morally valenced. To see a fellow creature as conscious (or sentient) often goes hand-in-hand with the sense that we should behave decently towards it (Singer, 1975; Nussbaum, 2023). To see an AI system as conscious would be to admit it into this fellowship of conscious beings, and potentially to give it “moral standing”, which would be a serious matter (Metzinger, 2021; Ladak, 2023).
Unfortunately, the ascription of consciousness, except in the ordinary case of human beings, is a tricky business, and the less human-like the candidate, the more problematic the ascription (Shevlin, 2021; Mudrik et al., 2023; Bayne et al., 2024). [^2] Even in humans, there are difficult clinical cases (Monti et al., 2010). Yet philosophical intuition is inclined to grant the possibility that, within the space of possible minds – a space that encompasses not only humans and other animals, but also putative extraterrestrial life-forms as well as sufficiently advanced kinds of artificial intelligence (Sloman, 1984) – there exist conscious exotica, entities that are extremely different from anything found in (terrestrial) biology but that belong to the fellowship of conscious beings (Nagel, 1974; Shanahan, 2016). [^3] The philosophical intuition in question manifests most prominently in the tropes of science fiction, but this is no reason to sideline it (Rennick, 2021).
@@
The central concern of this paper is whether (virtually embodied) LLM-based conversational agents, despite being “mere” simulacra of human behaviour, could ever qualify as conscious exotica. [^4] Butlin et al. (2023) address a similar question using a very different methodology. See also (Chalmers, 2023a). Following the treatment in (Shanahan, 2016), which draws heavily on the later work of Wittgenstein, the paper foregrounds embodied interaction as a basis for the use of consciousness language.
@@ 2   Language Models and AI
The core component of a contemporary conversational agent, such as OpenAI’s ChatGPT, Google’s Gemini, or Anthropic’s Claude, is a large language model, such as GPT-4 (OpenAI, 2023), Claude 3 (Anthropic, 2024), or Gemini Ultra (Anil et al., 2023). The basic task of the LLM is to (probabilistically) generate continuations of sequences of words (tokens) (Vaswani et al., 2017; Shanahan, 2024). The LLM embedded in a typical conversational agent is the product of two steps. First, a base model is trained to perform next token prediction on a large corpus of textual data. Second, the base model is fine-tuned a) to be effective at following instructions in a dialogue setting, and b) to take account of feedback from humans raters with respect to toxicity, bias, and so on. The resulting LLM is embedded in a dialogue system that takes turns with the user to build up a conversation in the context of an initial dialogue/system prompt, not seen by the user, that sets the tone of the exchange.
@@
Today’s conversational agents are hardly without their limitations. They have a tendency to generate inaccurate, made-up information (a phenomenon often (mis-)termed “hallucination”), and their reasoning skills are poor. Nevertheless, the experience of interacting with them is sufficiently compelling, and their conversational capabilities are sufficiently close to human level, that the urge to speak of them in anthropomorphic terms is almost overwhelming. [^5] For example, see the following conversation that took place in March 2024 between the present author and Anthropic’s Claude 3 Opus, covering a wide range of topics, including consciousness, selfhood, Buddhism, multiversality, and hyperstition: https://www.doc.ic.ac.uk/~mpsha/conversation_with_claude_march_2024_1.pdf.
Anthropomorphising AI systems is sometimes harmless fun, if the user is in the know, and sometimes useful for explaining and predicting a system’s behaviour, as when we adopt what Dennett calls the intentional stance, the strategy of “interpreting the behavior of an entity … as if it were a rational agent” (Dennett, 2009). Anthropomorphism is problematic when it involves the misleading attribution of human properties to systems that lack those properties, giving rise to false expectations for how the system will behave.
@@ 3   Anthropomorphism and Role Play
The position adopted in this paper, following (Shanahan, 2024), is that we should be wary of taking too seriously talk of beliefs in simple LLM-based conversational agents, despite their impressive conversational skills, since they lack the means to “participate fully in the human language game of truth”. Specifically, a simple LLM-based conversational agent, according to the definition of “simple” in use here, cannot measure its words against external reality and update what it says accordingly, a capacity that is central to the concept of belief in its fullest sense.
While it may not be appropriate to think of simple conversational agents as literally having beliefs, they can still usefully be thought of as role-playing or simulating an entity with beliefs (Andreas, 2022; Janus, 2022; Shanahan et al., 2023). LLMs encode a great deal of human knowledge, and a suitably fine-tuned and prompted base model will effectively play the part of a helpful assistant in a turn-taking setting, answering factual questions (more or less accurately) as if it believed, and had good reasons to believe, its own answers. In general, the role-play framing allows us to use familiar folk-psychological terms to describe, explain, and predict the behaviour of LLM-based systems without falling into the trap of anthropomorphism.
@@
Now, to what extent can we extend this treatment of belief to other mental attributes, such as desires, goals, intentions, and, most pertinently, to consciousness? To a degree, the same trends that legitimise talk of belief in enhanced conversational agents apply to the concepts of goals and intentions. A conversational agent capable of deliberation and tool-use can play the part of an assistant that forms plans on behalf of the user, and sets about executing them by carrying out real-world actions such as making purchases, sending emails, and so on. In this narrow context, there is little to distinguish role play from authenticity. Real emails are sent, real items are purchased. Thus the agent fulfils its intentions and meets its goals, hopefully to the satisfaction of a real user.
@@ 4   Wittgenstein Versus Dualism
The aim of this section is to present a philosophical approach that can be applied to exotic candidates for admission to the fellowship of conscious beings, such as LLM-based conversational agents. First, though, some remarks in lieu of a definition are in order. Consciousness is a multi-faceted concept. In everyday conversation, we speak of wakefulness, awareness, attention, experience, sensation, feeling, emotion, and so on. The scientific and philosophical literature supplements common speech with a distinctive vocabulary of its own: perception, introspection, phenomenology, sentience, selfhood, higher-order states, mental imagery, inner speech, and so on. [^6] Relatedly, in the context of animal consciousness, Birch, et al. (2020) distinguish five aspects, or dimensions, of consciousness: perceptual richness, evaluative richness, integration at a time, integration across time, and self-consciousness. To speak generically of consciousness is to allude to this whole cloud of concepts, and an entity has the capacity for consciousness to the extent that this vocabulary is applicable to it. [^7] The influential distinction between phenomenal consciousness and access consciousness introduced by Block (1995) in a related context is set aside here, as it is inimical to the present philosophical project.
@@
To achieve this dissolution we have to take on board Wittgenstein’s overarching philosophical project, wherein his view of language plays a central role. According to this view, language is an inherently embodied and social phenomenon, an aspect of human collective activity. So rather than asking what a word means, we should instead ask how it is used, what its role is in everyday human affairs. This applies no less to tricky philosophical words, such as “consciousness” and its relatives, than it does to everyday words like “flower” or “hello”. Philosophical puzzles arise when “language goes on holiday”, [^9] Wittgenstein (1953) (henceforth PI) §38. when philosophically difficult words are taken far from their natural home in everyday life and used in peculiar ways to “bewitch our intelligence”. [^10] PI §109.
To undo the spell of dualism, in the case of words like “sensation”, “experience”, and “feeling”, is no easy matter, but in the private language remarks, Wittgenstein takes us through a series of steps with the aim of doing this. [^11] PI §§256–271. He shows that a word that purported to denote a purely private sensation could not have any possible use in our language. Only words whose correct usage can be adjudicated through what is public, by the community of language users, can have meaning. Wittgenstein is careful neither to deny nor to affirm the existence of the allegedly private, hidden thing, the “sensation itself”, so to speak. It is “not a something, but not a nothing either”. The conclusion, rather, is that “a nothing would serve as well as a something about which nothing can be said”. [^12] PI §304.
@@ 5.3 A society-wide conversation
There is no guarantee of consensus here. Disagreement and debate are part of the conversation. [^14] Humphrey, for example, takes issue with the current (near) consensus on octopus sentience on the grounds that “they are not natural psychologists, they do not regard each other as selves, nor do they care” (Humphrey, 2022, p.205). Nor, as the conversation progresses, is there any guarantee of eventual convergence. Even if we do begin to treat an exotic entity as a conscious being and to describe it in such terms, as time goes by and more is learned about it, either at the level of behaviour or at the mechanistic level, this tendency might fade. Perhaps we will decide, collectively, that the language of consciousness is not the right one after all. Perhaps a little more nuance will be required. Perhaps a whole new vocabulary will emerge. [^15] Putnam (1964) articulates a position not so far from the one advocated here: “[T]he question: Are robots conscious? calls for a decision, on our part, to treat robots as fellow members of our linguistic community, or not to so treat them. As long as we leave this decision unmade, the statement that robots [who use language] are conscious has no truth value.” (p.690).
@@ 6.1 Simple conversational agents
Even simple conversational agents built on large language models, agents that that do nothing more than engage in textual dialogue, can elicit the feeling of a presence at the other end of the conversation (Schwitzgebel, 2023). So it is perhaps unsurprising that some users who have spent lots of time interacting with these agents will start to think of them as fellow conscious beings and to speak of them in such terms (Tiku, 2022; colombatto2023folk; Guingrich and Graziano, 2023). [^16] If a human ascribes consciousness to an AI system even though they know that it is an artefact, then it passes the Garland Test, named after the 2015 film Ex Machina, written and directed by Alex Garland, in which the robot Ava is subjected to just such a test (Shanahan, 2016; Seth, 2021). The Garland Test is neutral about whether or not the artefact in question is really conscious (whatever that may mean). But how seriously should we take this?
According to the prescription of the previous section, for an exotic entity to qualify as as a candidate for the fellowship of conscious beings it must be possible to engineer an encounter with it, at least in principle if not in practice. However, it is not possible to engineer an encounter with a simple conversational agent, even in principle. This is because simple LLM-based conversational agents are not embodied; we cannot be with them in a shared world. [^17] There is a standard set of supposed counter-examples to this view that includes brains in vats, minds uploaded into computers, patients with locked-in syndrome, and perfectly normal people in sensory deprivation tanks. Each of these example is sometimes alleged to support the view that there can be consciousness without embodiment. However, in every case it is possible to engineer an encounter in the current sense. The brain in a vat can be re-connected to its body, the locked-in patient can be cured, the uploaded mind can be downloaded again, and the occupant of the sensory deprivation tank can be dragged back into the daylight.
@@
To use the language of consciousness in a disembodied setting is to stray impossibly far from all this, its original home. If some community of language users insists on doing so, then one of two things must hold. Either they have bent the language of consciousness so radically out of shape that it has detached from its original nexus of meaning. Or, to the extent they believe this is not the case, that “consciousness” means the same for them as it always did for everyone else, they are clinging to the (alluring) dualistic picture of consciousness as a metaphysical kind whose very nature (private, hidden) means that it does not require embodiment in a world shared with others. [^18] Chalmers, for example, defends the possibility of “pure thinkers”, conscious beings “that can think but that have never had the capacity to sense” (Chalmers, 2023b).
@@ 6.2 Physical embodiment
The LLM-based generative AI systems deployed by today’s major tech corporations, such as OpenAI’s ChatGPT, Anthropic’s Claude, and Google’s Gemini, are more than simple conversational agents. Although language models remain their core component, they are multi-modal, tool-using conversational agents. As well as text, they can take images as input and generate images as output, and they can make calls to external apps (tools) such as calendars, calculators, search engines, and Python interpreters. Additionally, the architecture of these agents can be elaborated in ways that bring them closer to humans and animals, equipping them with a persistent memory of their interactions with the user and the world, for example, and by giving them the ability to plan and to form explicit lists of tasks and sub-goals (Xi et al., 2023).
@@
Is there any way to extend an LLM-based conversational agent into an AI system with which we could engineer an encounter, and that would thereby qualify as a candidate for consciousness? Indeed there is. We could embody it. One possibility is embodiment in physical robot form (Driess et al., 2023; Brohan et al., 2023; Yoshida et al., 2024). An LLM with multi-modal, tool-using capabilities is half-way there. Robot actions become another sort of tool, while data from the robot’s sensors (including camera images) is assimilated into its multi-modal input space. The LLM can then be suitably fine-tuned and prompted to carry out tasks for a user or to follow their instructions.
@@
Biological brains evolved to enable animals to survive and reproduce in complex environments, and language evolved to serve those fundamental needs. [^19] For some authors, consciousness is similarly bound up with these biological fundamentals. Seth, for example, writes: “[A]ll of our experiences and perceptions stem from our nature as self-sustaining living machines that care about their own persistence” (2021, p.255), while Aru, et al. (2023) question whether it is possible to “abstract consciousness away from the organizational complexity that is inherent within living systems but strikingly absent from AI systems”. In nature, therefore, embodiment is given, and language is inherently grounded in interaction with the physical world. In a robot controlled by an LLM, all this is back-to-front. A pre-trained statistical model of human language is bolted post hoc onto a robot body with no fundamental biological needs. The exoticism of the language model itself is compounded accordingly; a robot controlled by an LLM that exhibited human-like behaviour would be an especially exotic artefact.